# Quantifying Emotional Arousal through Pupillary Response: A Novel Approach for Isolating the Luminosity Effect and Predicting Affective States

### v1

@@ II-A Development of the Luminosity Effect Prediction Model
We first developed the LEPM model to isolate emotional pupil responses and subtracted its predictions from the observed pupil sizes. The LEPM development process involved several steps. First, we created a model that predicted the luminosity of a computer screen. Second, we developed a model that predicted pupil size in a dark laboratory using monochrome images (red, green, blue, and gray) with varying luminosities, which were tested on monochrome and non-monochrome emotionally neutral images. Third, we validated the model in a well-lit laboratory environment.
II-A1 Computing the luminosity of a monochrome image on a computer screen in a dark laboratory
@@
where is the luminosity value, is a constant (scaling factor) that depends on the brightness/contrast settings of the monitor, and are the RGB intensity values of a monochrome image displayed on the screen [pansara2024towards]. The relationship between RGB intensity and luminosity values is non-linear for a commercial monitor. For example, considering a primary color, e.g., blue, the luminosity corresponding to an intensity value of 60% is not twice as high as that corresponding to a value of 30%. Instead, the increase in luminosity as a function of RGB intensity is smaller at lower intensity values and dramatically increases as the intensity approaches maximum [pansara2024towards]. Furthermore, as more colors are used, this non-linear relationship changes. Therefore, the function in Equation 1 cannot be easily described analytically; therefore, we tabulated its values in a look-up table to accurately calculate the luminosity of images or frames displayed on a computer screen. The input of our look-up table was an RGB value corresponding to three integers ranging from 0 to 100. Luminosity values were measured empirically. We created a set of images with different RGB values. The luminosity of each image was measured in a dark laboratory using an LX1010BS digital lux meter at a distance of 65 cm, following standard eye-tracking protocols to ensure controlled conditions for accurate gaze and pupil response measurements [pansara2024towards]. If we wanted to consider all values, we would have entries and would have to generate images. We sub-sampled this space to make the search computationally less expensive, producing 1330 images. When querying the table, the corresponding luminosity value was returned if the input matched a point. Otherwise, a weighted average of the eight nearest points was calculated, with weights inversely proportional to their distances. In the case of images composed of only one of the fundamental colors, instead of eight nearest neighbors, we used only two [pansara2024towards].
@@
We calculated the values in the lookup table using a Dell Precision M6500 (1920x1080) monitor set to 100% brightness and 100% contrast, which we will refer to as the “reference monitor” [pansara2024towards]. Different monitor models have different luminosity capabilities. However, when testing five monitors of different sizes and brands, we consistently observed that the non-linear relationship described in equation 1 remains substantially valid, except for the scale factor which must be recalculated every time the brightness and contrast settings are changed (see calibration procedure below). In particular, once the appropriate value of was found, the maximum error between the monitors was 3 lux [pansara2024towards], which is tolerable for our purposes. This suggests that, regardless of the manufacturer and model of the monitor, the relationship between the brightness values and RGB intensity is highly consistent between different monitors. The calibration procedure described below allowed us to compensate for the differences between the possible monitor settings and between different monitors, provided that the monitor settings did not change during the experiment. For the reference monitor, the value of was 1 [pansara2024towards].
@@
II-A2 Computing the luminosity of a non-monochrome image on a computer screen in a dark laboratory
@@
When an image or video frame is not monochrome, one can calculate the luminosity of each pixel and average the values obtained. However, we observed that the luminosity of an image is similar to that of a single-color image, with the average RGB value of the original image calculated by averaging the RGB values of all its pixels. To validate this observation, we selected 100 images: 50 taken from the Internet and 50 generated by assigning a random RGB value to each pixel. We then generated the corresponding monochrome images using the average RGB value of each original image. The luminosity of the original images differed from the luminosity of the corresponding monochrome images by 1.5 lux on average, with a maximum difference of 3 lux, tolerable for our purpose. For computational reasons, using the corresponding monochrome images is preferable, as it is much faster than calculating the luminosity for each pixel by querying our look-up table and then averaging [pansara2024towards].
II-A3 Predicting pupil size due to luminosity at different wavelengths
@@
The second step consisted of developing a model to predict the pupil size of a participant looking at a monochrome image on the screen111Throughout this study, we used an average of right and left pupil size as pupil size data.. This model incorporated the function , defined in the previous step, which transforms RGB values into luminosity. This luminosity function then served as an input to another function responsible for mapping luminosity to pupil size. Notably, this second function is highly nonlinear and exhibits a decreasing trend [zhang2019pupil].
For each primary color plus gray, the model can be described using the following equation:
@@
where is the pupil size, is part of the luminosity function that gives the predicted luminosity (see Equation 1), and , , , and are coefficients to be determined by fitting the model to the recorded data [pansara2024towards]. The parameters and also absorb the parameter in the Equation 1. Our goal was to assess the pupil response either to one of the fundamental wavelengths (red, green, blue) or to their uniform combination (gray).
@@
To fit the model shown in Equation 2, for each of the four colors, we performed an initial estimation of the coefficients using data recorded from ten participants and then devised a method to recalibrate the coefficients , , , and to new experimental participants. We sequentially presented 101 monochrome frames (video) of each of the four colors to the 10 participants for the initial estimation. We started from the lowest intensity (0%, 0%, 0%) to the highest intensity (e.g., for red 100%, 0%, 0% and for gray 100%, 100%, 100%), increasing the brightness by 1% every second. Each of the four videos lasted 101 seconds and was presented full-screen on the reference monitor (1920x1080). We measured pupil size for each video frame at a data rate of approximately222The data rate of the pupil size measured with the Tobii Pro Nano and iMotions software is variable. 60 Hz and at a distance of 65 cm between the screen and the participant, using a Tobii Nano Pro eye tracker and the iMotions software (https://imotions.com/) in a dark (unlit) laboratory. At the same time, we measured luminosity during each video frame using the lux meter [pansara2024towards]. When an eye blink occurred (indicated by a value of -1 in the dataset output of the iMotions software), it was replaced with an estimated pupil size, which was calculated as the median pupil size for that specific frame. For each RGB intensity and participant, we considered the average pupil size recorded during the video frame presentation corresponding to that particular luminosity. Subsequently, we calculated the average pupil size data across all participants for each luminosity level. Figure 1 displays the average pupil size (aggregated across participants) plotted against luminosity for each color (scattered points). The dotted lines represent the fitted models333We did not need 101 measurements to fit a model with four coefficients, but using 101 data points allowed us to evaluate whether the model of the Equation 2 was the most appropriate.. As shown, pupil size diminished exponentially with rising luminosity [zhang2019pupil, pansara2024towards]. Figure 1 shows that the chosen model appropriately describes the data (R2 0.98).
@@
Table I shows the coefficient values of the model presented in Figure 1 [pansara2024towards]. These were then re-calibrated for each new participant. Indeed, it is essential to redetermine the coefficients before starting an experiment and tailoring them to the individual participant. To achieve this, each participant completed a calibration procedure before any experiment by watching a video of 27 monochrome frames [pansara2024towards]. The frames consisted of four colors –red, blue, green, and gray– each displayed for 4 seconds, resulting in a total duration of 108 seconds. These images were generated using all possible combinations of three RGB intensity levels 0%, 50%, and 100%, covering a range from the darkest (0%, 0%, 0%) to the brightest (100%, 100%, 100%).
@@
II-A4 Two different approaches to predict the pupil size of a participant looking at a non-primary color monochrome image in a dark laboratory
@@
We pursued this goal with two approaches. The first approach, which we called “gray-based,” considered the pupil variation dependent only on the total luminosity of the image and not on the particular color. On the other hand, the second approach, called “color-based,” considered the luminosity of each primary color differently in the image. We expected that the first approach would work best for images where the three primary colors were fairly balanced, while the second approach would work best for images where one of the three primary colors was prevalent.
The “gray-based” approach involved calculating the image luminosity on the screen as a weighted average of the luminosities of the eight nearest images in the look-up table, as explained above. We then calculated the relative pupil size as if the image were equally composed of the three primary colors (according to the definition of gray used in this paper). In practice, we used the fourth equation of the set of Equations 2. To evaluate the “gray-based” approach and the calibration procedure, we tested 18 participants across five different monitors with varying resolutions, ranging from a minimum of to a maximum of . The test measurements included 18 () images from the calibration video and nine additional test images.
@@
where is predicted pupil size. The weights of each contribution to in Equation 3 are equal to the contribution of each fundamental color in the initial image [pansara2024towards]. For example, for a picture with RGB values , we have:
@@
II-A5 Combining the “gray-based” and the “color-based” approaches
@@
II-A6 Testing the Luminosity Effect Prediction Model in a well-lit laboratory and with non-monochrome images
@@
The results obtained with the non-monochrome images were not as promising as those obtained with monochrome images. In particular, images that contained a very bright region within a dark figure yielded much worse results than monochrome images. This was because an image could be dark on average, but the participant looked at the bright region, potentially inducing a more pronounced pupillary constriction than that elicited by an image with its average luminosity level. Therefore, it was necessary to determine where participants were looking through the recorded eye-tracking data. Specifically, we analyzed a region with a 300-pixel radius centered on where the gaze was directed at each instant. If the average luminosity of that specific region exceeded that of the average luminosity of the entire image, we used that region average instead. This method provided a more accurate representation of the subjective luminance and led to improved results, which we present in the Results Section III-A.
II-B Development of Arousal Detection Model
@@
The second part of the study aimed to demonstrate that the pupil size value discounted from luminosity was immediately usable for detecting emotional arousal levels without the need for complex pre-processing, advanced machine learning, or even deep learning models.
@@
We recruited 55 participants who, after the calibration procedure described above, watched 32 video clips with emotional content while pupil size was recorded. After each video clip, participants self-reported the experienced emotions on a scale of 0 to 9 for 12 emotions (more details in Data Collection part II-B1), which was used as the target value to predict self-reported arousal (more details in Arousal Ground Truth part II-B2).
II-B1 Data Collection
@@
The categorization of emotions has long been debated, with two dominant approaches: basic emotion theory, which defines emotions as distinct categories, and multi-dimensional theory, which conceptualizes emotions as points along interconnected scales [zhang2020emotion]. This ongoing debate highlights the complexity of understanding human emotions. Our study adopted the multi-dimensional theory, utilizing Russell’s widely recognized circumplex model [russell1980circumplex], which maps emotions along two key dimensions: valence (pleasantness) and arousal (activation level). Valence represents emotions ranging from unpleasant to pleasant, while arousal captures the intensity of emotional states, from calm to highly energized. These two dimensions are widely recognized in affective computing and cognitive research as a robust framework for modeling emotional experiences [kim2016identifying, ramaswamy2024multimodal] and have the advantage of providing a comprehensive yet efficient representation of complex emotional states.
@@
We selected 32 video clips with emotional content. These were classified into four distinct groups based on Russell’s circumplex model [russell1980circumplex]: high arousal with positive valence, high arousal with negative valence, low arousal with positive valence, and low arousal with negative valence (this classification was tested through a pilot study with 37 participants).
@@
As part of the participant recruitment process, we conducted a comprehensive screening procedure to identify potential indicators of conditions such as alexithymia, anxiety, depression, and personality disorders, which are known to influence emotional processing and responses [da2018alexithymia, gray2021emotion, kret2015emotion]. 130 participants were initially screened, with 55 meeting the inclusion criteria, by which eligible participants must be 18–65 years old, with sufficient English language proficiency to sign consent and follow questionnaires, but without severe psychiatric conditions affecting emotional recognition (e.g., schizophrenia, bipolar disorder, autism; we accepted mild anxiety, depression, or alexithymia as measure by GAD-7, PHQ-9, and TAS-20), without intellectual disabilities, significant cognitive impairments, or recent neurological disorders (e.g., epilepsy, stroke, TBI in the past year). The final sample enrolled in the study comprised 26 male and 29 female participants, aged 18 to 65 years (mean: 27.79 10.360666In this paper, after the symbol , we always indicate the standard deviation and never the standard error of the mean.).
@@
All participants provided informed consent before participation, and the study was approved by the University of Essex Ethics Committee (reference number: ETH2425-0176) as part of a more extensive study that included recording pupil size, facial expression data, galvanic skin response, photoplethysmogram, and electroencephalogram. However, we only report on measurement and data processing related to pupil size. The iMotions software (https://imotions.com/) [imotionsPupillometry101] managed the experimental protocol and facilitated efficient data acquisition.
@@
The experiment was conducted in a well-lit laboratory environment (see Figure 4). At the beginning of the experiment, and before displaying the video clips, participants were shown the calibration video described in section II-A. Participants were instructed to focus on the center of the screen during the calibration procedure.
@@
Following calibration, the main experiment started. Each trial began with a 5-second gray screen displaying a central cross to neutralize previous emotional states, followed by an emotional video clip. We used 32 clips with a duration variable between 10 and 100 seconds, with different durations distributed uniformly across valence and arousal values, and shown in a randomized order. To determine the emotional arousal, we only considered the salient intervals from each video (see below). After each clip, participants completed a questionnaire (extended from [kim2016identifying] to include a broader range of emotions) to assess 12 emotional states: positive, negative, happy, calm, content, amused, excited, angry, sad, disgusted, fearful, and bored. Participants rated their emotional responses on a scale from 0 (no emotion) to 9 (high emotion).
@@
The final sample comprised 47 participants after excluding those based on three criteria: poor Tobii calibration that affected gaze and pupil size measurements, insufficient focus (e.g., participants consistently not looking at the screen or appearing to be asleep), and inconsistent responses, such as identical ratings for all questions. The excluded participants also showed low inter-rater reliability scores (Krippendorff’s alpha 0.4) as their answers were highly inconsistent with the other participants’ answers. The pupil size data from these remaining participants were then pre-processed as explained in section II-A.
II-B2 Arousal Ground Truth

### v2

@@ 2.1 Development of the Luminosity Effect Prediction Model
The LEPM development process involved several steps. First, we created a model that predicted the luminosity of a computer screen. Second, we developed a model that predicted pupil size in a dark laboratory using monochrome images (red, green, blue, and gray) with varying luminosities, which were tested on monochrome and non-monochrome emotionally neutral images. Third, we validated the model in normal laboratory conditions laboratory environment.
2.1.1 Computing the luminosity of a monochrome image on a computer screen in a dark laboratory
@@
where is the luminosity value, is a constant (scaling factor) that depends on the brightness/contrast settings of the monitor, and are the RGB intensity values of a monochrome image displayed on the screen Pansara et al. (2024). The relationship between RGB intensity and luminosity values is non-linear for a commercial monitor. For example, considering a primary color, e.g., blue, the luminosity corresponding to an intensity value of 60% is not twice as high as that corresponding to a value of 30%. Instead, the increase in luminosity as a function of RGB intensity is smaller at lower intensity values and dramatically increases as the intensity approaches maximum Pansara et al. (2024). Furthermore, as more colors are used, this non-linear relationship changes. Therefore, the function in Equation 1 cannot be easily described analytically; we tabulated its values in a look-up table to accurately calculate the luminosity of images or frames displayed on a computer screen. The input of our look-up table was an RGB value corresponding to three integers ranging from 0 to 100. Luminosity values were measured empirically. We created a set of images with different RGB values. The luminosity of each image was measured in a dark laboratory using an LX1010BS digital lux meter at a distance of 65 cm, following standard eye-tracking protocols to ensure controlled conditions for accurate gaze and pupil response measurements Pansara et al. (2024). If we wanted to consider all values, we would have entries and would have to generate images. We sub-sampled this space to make the search computationally less expensive, producing 1330 images. When querying the table, the corresponding luminosity value was returned if the input matched a point. Otherwise, a weighted average of the eight nearest points was calculated, with weights inversely proportional to their distances. In the case of images composed of only one of the fundamental colors, instead of eight nearest neighbors, we used only two Pansara et al. (2024).
@@
We calculated the values in the lookup table using a Dell Precision M6500 24-inch monitor (0.34 steradians at 65 cm) set to 100% brightness and 100% contrast, which we will refer to as the “reference monitor” Pansara et al. (2024). Different monitor models have different luminosity capabilities. However, when testing five monitors of different sizes and brands, we consistently observed that the non-linear relationship described in equation 1 remains substantially valid, except for the scale factor which must be recalculated every time the brightness and contrast settings are changed (see calibration procedure below). In particular, once the appropriate value of was found, the maximum error between the monitors was 3 lux Pansara et al. (2024), which is tolerable for our purposes. This suggests that, regardless of the manufacturer and model of the monitor, the relationship between the brightness values and RGB intensity is highly consistent between different monitors. The calibration procedure described below allowed us to compensate for the differences between the possible monitor settings and between different monitors, provided that the monitor settings did not change during the experiment. For the reference monitor, the value of was 1 Pansara et al. (2024).
@@
2.1.2 Computing the luminosity of a non-monochrome image on a computer screen in a dark laboratory
@@
When an image or video frame is not monochrome, one can calculate the luminosity of each pixel and average the values obtained. However, we observed that the luminosity of an image is similar to that of a single-color image, with the average RGB value of the original image calculated by averaging the RGB values of all its pixels. To validate this observation, we selected 100 images: 50 taken from the Internet and 50 generated by assigning a random RGB value to each pixel. We then generated the corresponding monochrome images using the average RGB value of each original image. The luminosity of the original images differed from the luminosity of the corresponding monochrome images by 1.5 lux on average, with a maximum difference of 3 lux, tolerable for our purpose. For computational reasons, using the corresponding monochrome images is preferable, as it is much faster than calculating the luminosity for each pixel by querying our look-up table and then averaging Pansara et al. (2024).
2.1.3 Predicting pupil size due to luminosity at different wavelengths
@@
The second step consisted of developing a model to predict the pupil size of a participant looking at a monochrome image on the screen111Throughout this study, we used an average of right and left pupil size as pupil size data.. This model incorporated the function , defined in the previous step, which transforms RGB values into luminosity. This luminosity function then served as an input to another function responsible for mapping luminosity to pupil size. For the choice of this second function, we were inspired by the works mentioned in the introduction and in particular by the work of Pamplona et al. Pamplona et al. (2009) and the unified formula of Watson and Yellott Watson and Yellott (2012), readapted by Fanourakis and Chanel Fanourakis and Chanel (2022). All these works present a non-linear decreasing function that can be well approximated by an exponential, as also confirmed by Zhang et al. Zhang et al. (2019). For each primary color plus gray, our model can be described using the following equations:
@@
where is the pupil size, is part of the luminosity function that gives the predicted luminosity (see Equation 1), and , , , and are coefficients to be determined by fitting the model to the recorded data Pansara et al. (2024). The parameters and also absorb the parameter in the Equation 1. Our goal was to assess the pupil response either to one of the fundamental wavelengths (red, green, blue) or to their uniform combination (gray).
@@
To fit the model shown in Equation 4, for each of the four colors, we performed an initial estimation of the coefficients using data recorded from ten participants and then devised a method to recalibrate the coefficients , , , and to new experimental participants. We sequentially presented 101 monochrome frames (video) of each of the four colors to the 10 participants for the initial estimation. We started from the lowest intensity (0%, 0%, 0%) to the highest intensity (e.g., for red 100%, 0%, 0% and for gray 100%, 100%, 100%), increasing the brightness by 1% every second. Each of the four videos lasted 101 seconds and was presented full-screen on the reference monitor. We measured pupil size for each video frame at a data rate of approximately222The data rate of the pupil size measured with the Tobii Pro Nano and iMotions software is variable. 60 Hz and at a distance of 65 cm between the screen and the participant, using a Tobii Nano Pro eye tracker and the iMotions software (https://imotions.com/) in a dark (unlit) laboratory. At the same time, we measured luminosity during each video frame using the lux meter Pansara et al. (2024). When an eye blink occurred (indicated by a value of -1 in the dataset output of the iMotions software), it was replaced with an estimated pupil size, which was calculated as the median pupil size for that specific frame. For each RGB intensity and participant, we considered the average pupil size recorded during the video frame presentation corresponding to that particular luminosity. Subsequently, we calculated the average pupil size data across all participants for each luminosity level. Figure 1 displays the average pupil size (aggregated across participants) plotted against luminosity for each color (scattered points). The dotted lines represent the fitted models333We did not need 101 measurements to fit a model with four coefficients, but using 101 data points allowed us to evaluate whether the model of the Equation 4 was the most appropriate.. As shown, pupil size diminished exponentially with rising luminosity Zhang et al. (2019); Pansara et al. (2024). Figure 1 shows that the chosen model appropriately describes the data ( 0.98).
@@
Table 1 shows the coefficient values of the model presented in Figure 1 Pansara et al. (2024). These were then re-calibrated for each new participant. Indeed, it is essential to redetermine the coefficients before starting an experiment and tailoring them to the individual participant. To achieve this, each participant completed a calibration procedure before any experiment by watching a video of 27 monochrome frames Pansara et al. (2024). The frames consisted of four colors –red, blue, green, and gray– each displayed for 4 seconds, resulting in a total duration of 108 seconds. These images were generated using all possible combinations of three RGB intensity levels 0%, 50%, and 100%, covering a range from the darkest (0%, 0%, 0%) to the brightest (100%, 100%, 100%).
@@
2.1.4 Two different approaches to predict the pupil size of a participant looking at a non-primary color monochrome image in a dark laboratory
@@
We achieved this with two approaches. The first approach, which we called “gray-based,” considered the pupil variation dependent only on the total luminosity of the image and not on the particular color. On the other hand, the second approach, called “color-based,” considered the luminosity of each primary color differently in the image. We expected that the first approach would work best for images where the three primary colors were fairly balanced, while the second approach would work best for images where one of the three primary colors was prevalent.
The “gray-based” approach involved calculating the image luminosity on the screen as a weighted average of the luminosities of the eight nearest images in the look-up table, as explained above. We then calculated the relative pupil size as if the image were equally composed of the three primary colors (according to the definition of gray used in this paper). In practice, we used the fourth equation of the set of Equations 4. To evaluate the “gray-based” approach and the calibration procedure, we tested 18 participants across five different monitors with varying resolutions, ranging from a minimum of 24 inches to a maximum of 49 inches (from 0.34 to 0.68 steradians). The test measurements included 18 () images from the calibration video and nine additional test images.
@@
where is predicted pupil size. The weights of each contribution to in Equation 5 are equal to the contribution of each fundamental color in the initial image Pansara et al. (2024). For example, for a picture with RGB values , we have:
@@
2.1.5 Combining the “gray-based” and the “color-based” approaches
@@
2.1.6 Testing the Luminosity Effect Prediction Model in a normal laboratory conditions and with non-monochrome images
@@
The results obtained with the non-monochrome images were not as promising as those obtained with
@@
monochrome images. In particular, images that contained a very bright region within a dark figure yielded much worse results than monochrome images. This was because an image could be dark on average, but the participant looked at the bright region, potentially inducing a more pronounced pupillary constriction than that elicited by an image with its average luminosity level. Therefore, it was necessary to determine where participants were looking through the recorded eye-tracking data. Specifically, we analyzed a region with a 300-pixel radius centered on where the gaze was directed at each instant. If the average luminosity of that specific region exceeded that of the average luminosity of the entire image, we used that region average instead. This method provided a more accurate representation of the subjective luminance and led to improved results, which we present in the Results Section 3.1.
2.2 Development of Arousal Detection Model
@@
The second phase of the study aimed to demonstrate that the pupil size value discounted from luminosity was immediately usable for detecting emotional arousal levels without the need for complex pre-processing, advanced machine learning, or even deep learning models.
@@
We recruited 55 participants who, after the calibration procedure described above, watched 32 video clips with emotional content while pupil size was recorded. In order to facilitate a clear validation of the isolated pupil residual, the selected video clips were relatively short (with a duration variable between 10 and 100 second) and each eliciting a single primary emotional state (more details below). After each video clip, participants self-reported the experienced emotions on a scale of 0 to 9 for 12 emotions (more details in Data Collection part 2.2.1), which was used as the target value to predict self-reported arousal (more details in Arousal Ground Truth part 2.2.2).
2.2.1 Data Collection
@@
The categorization of emotions has long been debated, with two dominant approaches: basic emotion theory, which defines emotions as distinct categories, and multi-dimensional theory, which conceptualizes emotions as points along interconnected scales Zhang et al. (2020). This ongoing debate highlights the complexity of understanding human emotions. Our study adopted the multi-dimensional theory, utilizing Russell’s widely recognized circumplex
model Russell (1980), which maps emotions along two key dimensions: valence (pleasantness) and arousal (activation level). Valence represents emotions ranging from unpleasant to pleasant, while arousal captures the intensity of emotional states, from calm to highly energized. These two dimensions are widely recognized in affective computing and cognitive research as a robust framework for modeling emotional experiences Kim et al. (2016); Ramaswamy and Palaniswamy (2024) and have the advantage of providing a comprehensive yet efficient representation of complex emotional states.
@@
We selected 32 video clips with emotional content. These were classified into four distinct groups (8 videos per group) based on Russell’s circumplex model Russell (1980): high arousal with positive valence, high arousal with negative valence, low arousal with positive valence, and low arousal with negative valence (this classification was tested through a pilot study with 37 participants). The duration of clips varied between 10 and 100 seconds, with different durations distributed uniformly across valence and arousal values and shown in a randomized order.
@@
As part of the participant recruitment process, we conducted a comprehensive screening procedure to identify potential indicators of conditions such as alexithymia, anxiety, depression, and personality disorders, which are known to influence emotional processing and responses Da Silva et al. (2018); Gray et al. (2021); Kret and Ploeger (2015). We initially screened 130 participants, with 55 meeting the inclusion criteria. Eligible participants were required to be between 18 and 65 years of age and to have sufficient English proficiency to provide informed consent and complete the questionnaires. Exclusion criteria included severe psychiatric conditions that could affect emotional recognition (e.g., schizophrenia, bipolar disorder, or autism spectrum disorder); however, mild anxiety, depression, or alexithymia were permitted, as assessed by the GAD-7, PHQ-9, and TAS-20. Additional exclusion criteria were intellectual disability, significant cognitive impairment, or recent neurological disorders (e.g., epilepsy, stroke, or traumatic brain injury within the past year). Eligibility was assessed using online Qualtrics survey questionnaires. The final sample enrolled in the study comprised 26 male and 29 female participants, aged 18 to 65 years (mean: 27.79 10.360666In this paper, after the symbol , we always indicate the standard deviation and never the standard error of the mean.).
@@
All participants provided informed consent before participation, and the study was approved by the University of Essex Ethics Committee (reference number: ETH2425-0176) as part of a more extensive study that included recording pupil size, facial expression data, galvanic skin response, photoplethysmogram, and electroencephalogram. Here, we only report on measurement and data processing related to pupil size. The iMotions software (https://imotions.com/) iMotions managed the experimental protocol and facilitated efficient data acquisition.
@@
The experiment was conducted in a laboratory environment under normal lighting conditions, i.e., natural light during the day and in good weather, artificial light otherwise. (see Figure 4). At the beginning of the experiment, and before displaying the video clips, participants were shown the calibration video described in section 2.1. Participants were instructed to focus on the center of the screen during the calibration procedure.
@@
Following calibration, the main experiment started. Each trial began with a 5-second gray screen displaying a central cross to neutralize previous emotional states, followed by an emotional video clip. To determine the emotional arousal, we only considered the salient intervals from each video (see below). After each clip, participants completed a questionnaire (extended from Kim et al. (2016) to include a broader range of emotions) to assess 12 emotional states: positive, negative, happy, calm, content, amused, excited, angry, sad, disgusted, fearful, and bored. Participants rated their emotional responses on a scale from 0 (no emotion) to 9 (high emotion).
@@
The final sample comprised 47 participants after excluding those based on three criteria: poor Tobii calibration that affected gaze and pupil size measurements, insufficient focus (e.g., participants consistently not looking at the screen or appearing to be asleep), and inconsistent responses, such as identical ratings for all questions. The excluded participants also showed low inter-rater reliability scores (Krippendorff’s alpha 0.4) as their answers were highly inconsistent with the other participants’ answers. The pupil size data from these remaining participants were then pre-processed as explained in section 2.1.
2.2.2 Arousal Ground Truth